OpenStudy (anonymous):

Solve for x. x/5 - x/6 = 1/3

OpenStudy (omarbirjas):

Hello and welcome to Openstudy. I can assist you in the answering of this question. This is what you have correct? \[\\\ \frac{ x }{ 5 }-\frac{ x }{ 6}=\frac{ 1 }{ 3}\]

OpenStudy (omarbirjas):

If you are not going to be available for help I will leave so that I can tend to other OS users needs. Please tag me when you come back.